In a singkil dance, two men are carrying a




princess on a bamboo pole that is 5.0 m long and weighs 200N. If the princess weighs 450N and sits 1.5 m on0 right end, how much each man support?


1
Expert's answer
2022-04-13T09:21:46-0400

"F=200+450=650~N,"

"F_1=P_1\frac{l_1}2+P_2l_2=150~N,"

"F_2=F-F_1=500~N."